OpenStudy (anonymous):

a/2=(a-4)/2 what does a equal

OpenStudy (anonymous):

to start, break up the fraction on the right... a/2=a/2-4/2 move variables... 0=-2, no solution. Are you sure thats the correct equation?

OpenStudy (anonymous):